OpenStudy (precal):

Need to find the second derivative of x^4+y^4=64

myininaya (myininaya):

\[(x^4)'=4x^3 ; (y^4)=4y^3y' ; (64)'=0\]

myininaya (myininaya):

\[4x^3+4y^3y'=0\] solve for y' \[y'=\frac{-x^3}{y^3}\] Now apply the quotient rule!

myininaya (myininaya):

i will let you try

OpenStudy (precal):

ok give me a moment to type it out or draw it out

myininaya (myininaya):

yes draw and i will correct errors :)

OpenStudy (precal):